The weight density of a fluid is determined to be 9683 N/m3.

Determine the relative density of this fluid.

Report your answer to three decimal places.
Answer #1

The given Weight density is 9683 N/m³

Weight density is product of (density (kg/m³ )× (g, acceleration due to gravity,m/s²)

So N= (kg.m/s²)

Now specific gravity is ratio of density of fluid to the density of water.

Mass density of fluid= (Weight density/g) (Kgm/s²/m³)/(m/s²)= kg/m³

=9683/9.81= 987.0454 kg/m³

And density of water is 1000 kg/m³

So specific gravity=(987.0454/1000)=0.987054

This can be also done if we just devide the weight density of fluid by the weight density of water which is 9.81×1000

Specific gravity=9683/9810= 0.987054

Give your feedback